# Zhuo Zhang College of Computer Science and Technology

### v2

@@ 2.3.3 Application-Specific Recommendations
Example: Yin et al. (Yin2024) demonstrated 1006 cardiac geometries with ¡2% error in electrophysiology simulations.
@@ 3.1.1 Incompressible Navier-Stokes Equations
1. Cylinder Flow Benchmark (Jin2021): Jin et al. developed NSFnets (Navier-Stokes Flow nets) combining residual networks with physics constraints for flow around circular cylinders parameterized by Reynolds number and cylinder diameter.
@@
4. Advanced Fluid Applications: The field has matured with production-grade applications. Qiu et al. (Qiu2025DNS) performed direct numerical simulations of 3D two-phase flow using physics-informed neural networks with distributed parallel training on Journal of Fluid Mechanics 2025, handling millions of degrees of freedom for bubble dynamics and interfacial phenomena. Li et al. (Li2024FV; Li2025FEM) developed finite volume-informed neural networks (Physics of Fluids 2024) and finite element-informed networks (JCP 2025) for unsteady incompressible flows, achieving data-free approaches that rely purely on physics residuals without any simulation data. Zou et al. (Zou2025FDIGN) integrated finite-difference schemes into graph networks in Physics of Fluids 2025 for flows on block-structured grids, achieving state-of-the-art accuracy for complex industrial geometries.
2. FNO for Kolmogorov Flow (Li2021ICLR): Li et al. applied Fourier Neural Operator to 2D forced turbulence, demonstrating unprecedented efficiency for parametric turbulent flow.
@@
3. Cavity Flow with DeepONet (Dong2022): Dong et al. addressed lid-driven cavity flow—a canonical benchmark in CFD—using DeepONet with multi-fidelity (Penwarden2023; Howard2023; Peherstorfer2018) data fusion.
@@
Results: Relative error: $8\%$ using multi-fidelity (Penwarden2023; Howard2023; Peherstorfer2018) approach vs $15\%$ with high-fidelity-only training on same budget, Captured critical Re transition where secondary vortices appear, and Inference: 0.1s per parameter configuration
@@ 3.1.2 Compressible Flows and Shocks
Recent Progress: Several approaches address shock capturing: Conservative formulation PINNs (Jagtap2020): Enforce conservation laws in integral form to handle weak solutions, Shock-capturing neural operators (List2023): Augment FNO with WENO-inspired local refinement near discontinuities, and Hyperbolic PINNs (Moseley2023): Domain decomposition along characteristic lines
[…]
@@ 3.1.3 Parametric Shape Optimization
State-of-the-Art: Li et al. (Li2023AeroDesign) demonstrated: Training: 30,000 airfoils from UIUC database, CFD simulations at $\text{Re}=10^{6}$, Architecture: Geo-FNO handling varying geometry through diffeomorphic mapping, Optimization: Evolutionary algorithm with 10,000 neural operator evaluations (30 seconds total) vs infeasible with CFD, and Result: Discovered novel airfoil design (Sekar2019)s with $L/D$ improvements of 8-12% over baseline
@@ 3.2.1 Linear Elasticity
1. Parameter Identification with PINNs (Haghighat2021): Haghighat et al. developed a PINN framework for identifying spatially-varying material properties from displacement measurements.
@@
2. Operator Learning for Parametric Structures (Goswami2022): Goswami et al. applied DeepONet to bridge structures parameterized by loading patterns and geometric variations.
@@ 3.2.2 Nonlinear Mechanics and Plasticity
Neural Constitutive Modeling: Recent work (Masi2021; Thakolkaran2022) learns constitutive relations directly from data:
@@
Application - Elastomeric Materials: Thakolkaran et al. (Thakolkaran2022) trained neural operators to predict large-deformation response of elastomers across parameter space of Ogden model coefficients. Achieved $2\%$ error in stress-strain curves for complex multiaxial loading, enabling rapid virtual testing for material design.
@@ 3.2.3 Topology Optimization
Zhu et al. (Zhu2023PhaseField) introduced Phase-Field DeepONet using energy-based loss functions for pattern formation, enabling fast Allen-Cahn and Cahn-Hilliard simulations. Lee et al. (Lee2025FEONO) developed FE Operator Networks for high-dimensional elasticity (d=50 parameters), achieving 60% error reduction.
@@
Neural Approach: Chandrasekhar and Suresh (Sosnovik2019) (Chandrasekhar2021) trained neural operators to map loading patterns to optimal topologies:
@@ 3.3.1 Parametric Heat Conduction
Industrial Application - Electronic Cooling: Cai et al. (Cai2021Heat) addressed thermal management of chip packages with varying power distributions and heat sink geometries.
@@ 3.3.2 Thermal Property Identification
PINN Approach: Anantha Padmanabha et al. (Anantha2021) developed Bayesian PINNs: Observations: Temperature at 50 locations via infrared thermography, Unknown: 2D conductivity field $k(x,y)$ with suspected discontinuities (material interfaces), and Method: PINN with total variation regularization to preserve sharp interfaces
@@ 3.3.3 Conjugate Heat Transfer
State-of-the-Art: Penwarden et al. (Penwarden2023) developed multi-fidelity (Penwarden2023; Howard2023; Peherstorfer2018) neural operators for conjugate heat transfer: Low fidelity: Simplified 1D thermal resistance models (cheap), High fidelity: Coupled CFD-conduction simulations (expensive), and Neural fusion: Multi-fidelity DeepONet learns correction from low to high fidelity
@@ 3.4.1 Parametric Maxwell’s Equations
Time-harmonic Maxwell equations (Chen2020Maxwell; Wiecha2021) in frequency domain:
@@
Application - Metamaterial Design: Chen et al. (Chen2020Maxwell) used neural operators for inverse design (Lu2021Adaptive; Meng2022) of electromagnetic metamaterials.
@@ 3.4.2 Acoustic Wave Propagation
The acoustic wave (Jin2022Acoustic) equation in heterogeneous media:
@@
Seismic Imaging Application: Moseley et al. (Moseley2023Seismic) addressed full-waveform inversion for subsurface velocity model estimation.
@@ 3.5 Cross-Domain Insights and Application Maturity
Fluid Dynamics Re, Mach, geometry 1-10 FNO, DeepONet, PINO $10^{3}$-$10^{5}\times$ (Li2021ICLR; Jin2021)
@@
Industrial Adoption Status: Production deployment: Weather forecasting (FourCastNet (Pathak2022; Kurth2023)), some engineering design optimization, Pilot projects: Digital twins, real-time control systems, medical imaging, and Research stage: Nuclear fusion, climate prediction, drug discovery
@@ 4.1.1 The Parametric Solution Manifold
Kolmogorov n-Width: A classical measure of manifold complexity from reduced-order modeling theory (Binev2011; Berkooz1993; Hesthaven2016; Peherstorfer2016) is the Kolmogorov $n$-width:
@@
For elliptic PDEs with smooth, affine parameter dependence, $d_{n}(\mathcal{M})\sim e^{-\alpha n}$ (exponential decay) justifies reduced basis methods (Binev2011). However, transport-dominated or hyperbolic problems may have $d_{n}\sim n^{-\beta}$ (polynomial decay), requiring larger $n$ for adequate approximation.
@@ 4.1.2 Universal Approximation Theorems for Operators
Theorem 2 (Neural Operator Universal Approximation (Kovachki2021)).
@@
DeepONet: Lu et al. (Lu2021) proved that the branch-trunk decomposition equation 28 achieves universal approximation if: Branch network width $p\to\infty$, and Both branch and trunk networks are universal approximators
@@
FNO: Li et al. (Li2021ICLR) showed Fourier neural operators approximate operators by leveraging spectral properties. For periodic problems, the Fourier series representation provides natural expressivity. The critical assumption: solutions have sufficient spectral decay (smoothness in Fourier space).
@@
where $W$ is network width, $d$ is input dimension, and $\alpha$ is smoothness. The curse of dimensionality ($1/W^{\alpha/d}$) persists theoretically (Han2018; Beck2021; Khoo2021), though empirically neural operators perform better than this bound suggests.
@@ 4.1.3 Parametric PDE-Specific Analysis
Low-Dimensional Structure: Bhattacharya et al. (Bhattacharya2021) proved that if the solution manifold $\mathcal{M}$ has intrinsic dimension $d_{\text{eff}}\ll d$ (as measured by Kolmogorov width decay), then neural operators can achieve dimension-independent error bounds:
@@ 4.2.2 Sample Complexity Bounds
De Ryck and Mishra (DeRyck2022) derived sample complexity for PINNs approximating parametric elliptic PDEs:
@@
For neural operators, Lanthaler et al. (Lanthaler2022) proved:
@@
Empirical Findings: FNO: Interpolation error $\sim$1-2%, extrapolation error $\sim$10-20% (Li et al. (Li2021ICLR)), DeepONet: Similar interpolation/extrapolation gap, slightly better extrapolation when physics-informed, and PINN: Poor parametric generalization—essentially no extrapolation capability without retraining
@@ 4.2.4 Out-of-Distribution Detection
Approaches: Uncertainty quantification employs multiple complementary approaches. Bayesian neural operators compute predictive variance as an out-of-distribution indicator. Ensemble methods identify uncertain regions through high variance across ensemble members. Conformal prediction constructs prediction sets with guaranteed coverage (Staber2024), providing distribution-free uncertainty bounds.
Breakthrough - Conformal Prediction for PDEs: Staber et al. (Staber2024) developed distribution-free uncertainty quantification for neural operators:
@@ 4.4.1 Training Convergence for PINNs
Neural Tangent Kernel (NTK) Analysis: Wang et al. (Wang2022NTK), building on failure mode analyses (Krishnapriyan2021; Krishnapriyan2022; Markidis2021; Fuks2020), used NTK theory to analyze PINN training dynamics. For infinitely-wide networks in the NTK regime:
@@
This explains empirically observed training failures (Krishnapriyan2021).
@@ 4.4.2 Operator Learning Convergence
Lanthaler et al. (Lanthaler2022) proved that for appropriate architecture choices and sufficient training data:
@@ 4.4.3 Long-Time Stability
Recent Success: Brandstetter et al. (Brandstetter2022) demonstrated stable 1000-step rollouts for turbulent flows by enforcing energy and enstrophy conservation in FNO training.
@@ 5.1.2 Active Subspaces and Dimension Reduction
Many high-dimensional parameter spaces have low effective dimensionality. Active subspace methods (Constantine2015) identify important parameter directions.
@@
Case Study: Hu et al. (Hu2024) addressed a subsurface flow (Tang2021; Fuks2020) problem with $d=100,000$ uncertain permeability parameters (each grid cell). Active subspace analysis revealed $d_{\text{eff}}=5$ active directions explaining 95% of pressure variance. After reduction: Trained DeepONet on 5D active subspace, Achieved $3\%$ error in pressure predictions, and Total computation: 12 hours on single GPU (vs. years for full-space sampling)
@@ 5.1.4 Latent Variable Models
L-DeepONet: Kontolati et al. (Kontolati2024) developed latent-space DeepONet for high-dimensional parametric PDEs: Application: Structural dynamics with $d=10,000$ (discretized forcing field) (Rao2023; Rezaei2022; Samaniego2020), Latent dimension: $d_{z}=50$ (learned via variational autoencoder), and Results: 1-2 orders of magnitude speedup vs. standard DeepONet, $5\%$ accuracy loss
@@ 5.2.1 Bayesian Physics-Informed Neural Networks
B-PINNs (Yang2021) place priors over network weights:
@@
Application - Cardiovascular Flows: Arzani et al. (Arzani2021), along with related cardiovascular modeling studies (Kissas2020; Sahli2022; Yazdani2020), used B-PINNs to infer patient-specific blood viscosity and vessel compliance from sparse Doppler ultrasound data. Posterior distributions quantified parameter uncertainty, enabling risk assessment (e.g., probability of flow reversal).
@@ 5.2.2 Bayesian Neural Operators
Efficient Approximations: Laplace approximation: Gaussian approximation around MAP estimate, Concrete dropout: Approximate variational inference via dropout (Gal2016), and Ensemble methods: Train multiple operators with different initializations

### v3

@@ 2.3.3 Application-Specific Recommendations
Example: Yin et al. (Yin et al., 2024) demonstrated 1006 cardiac geometries with ¡2% error in electrophysiology simulations.
@@ 3.1.1 Incompressible Navier-Stokes Equations
1. Cylinder Flow Benchmark (Jin et al., 2021): Jin et al. developed NSFnets (Navier-Stokes Flow nets) combining residual networks with physics constraints for flow around circular cylinders parameterized by Reynolds number and cylinder diameter.
@@
4. Advanced Fluid Applications: The field has matured with production-grade applications. Qiu et al. (Qiu et al., 2025) performed direct numerical simulations of 3D two-phase flow using physics-informed neural networks with distributed parallel training on Journal of Fluid Mechanics 2025, handling millions of degrees of freedom for bubble dynamics and interfacial phenomena. Li et al. (Li et al., 2024a; 2025c) developed finite volume-informed neural networks (Physics of Fluids 2024) and finite element-informed networks (JCP 2025) for unsteady incompressible flows, achieving data-free approaches that rely purely on physics residuals without any simulation data. Zou et al. (Zou et al., 2025) integrated finite-difference schemes into graph networks in Physics of Fluids 2025 for flows on block-structured grids, achieving state-of-the-art accuracy for complex industrial geometries.
2. FNO for Kolmogorov Flow (Li et al., 2020a): Li et al. applied Fourier Neural Operator to 2D forced turbulence, demonstrating unprecedented efficiency for parametric turbulent flow.
@@
3. Cavity Flow with DeepONet (Dong and Li, 2021): Roohi et al. addressed lid-driven cavity flow—a canonical benchmark in CFD—using DeepONet with multi-fidelity (Penwarden et al., 2022; Lu et al., 2022b; Peherstorfer et al., 2018) data fusion.
@@
Results: Relative error: $8\%$ using multi-fidelity (Penwarden et al., 2022; Lu et al., 2022b; Peherstorfer et al., 2018) approach vs $15\%$ with high-fidelity-only training on same budget, Captured critical Re transition where secondary vortices appear, and Inference: 0.1s per parameter configuration
@@ 3.1.2 Compressible Flows and Shocks
Recent Progress: Several approaches address shock capturing: Conservative formulation PINNs (Jagtap and Karniadakis, 2020): Enforce conservation laws in integral form to handle weak solutions, Shock-capturing neural operators (List et al., 2022): Augment FNO with WENO-inspired local refinement near discontinuities, and Hyperbolic PINNs (Moseley et al., 2020b): Domain decomposition along characteristic lines
@@ 3.2.1 Linear Elasticity
1. Parameter Identification with PINNs (Haghighat et al., 2021): Haghighat et al. developed a PINN framework for identifying spatially-varying material properties from displacement measurements.
@@
2. Operator Learning for Parametric Structures (Goswami et al., 2022): Goswami et al. applied DeepONet to bridge structures parameterized by loading patterns and geometric variations.
@@ 3.2.2 Nonlinear Mechanics and Plasticity
Neural Constitutive Modeling: Recent work (Masi et al., 2021; Thakolkaran et al., 2022) learns constitutive relations directly from data:
@@
Application - Elastomeric Materials: Thakolkaran et al. (Thakolkaran et al., 2022) trained neural operators to predict large-deformation response of elastomers across parameter space of Ogden model coefficients. Achieved $2\%$ error in stress-strain curves for complex multiaxial loading, enabling rapid virtual testing for material design.
@@ 3.2.3 Topology Optimization
Zhu et al. (Li et al., 2023a) introduced Phase-Field DeepONet using energy-based loss functions for pattern formation, enabling fast Allen-Cahn and Cahn-Hilliard simulations. Lee et al. (Lee et al., 2025) developed FE Operator Networks for high-dimensional elasticity (d=50 parameters), achieving 60% error reduction.
@@
Neural Approach: Chandrasekhar and Suresh (Sosnovik and Oseledets, 2019) (Chandrasekhar and Suresh, 2021) trained neural operators to map loading patterns to optimal topologies:
@@ 3.3.1 Parametric Heat Conduction
Industrial Application - Electronic Cooling: Cai et al. (Cai et al., 2021) addressed thermal management of chip packages with varying power distributions and heat sink geometries.
@@ 3.3.2 Thermal Property Identification
PINN Approach: Anantha Padmanabha et al. (Padmanabha and Zabaras, 2021) developed Bayesian PINNs: Observations: Temperature at 50 locations via infrared thermography, Unknown: 2D conductivity field $k(x,y)$ with suspected discontinuities (material interfaces), and Method: PINN with total variation regularization to preserve sharp interfaces
@@ 3.3.3 Conjugate Heat Transfer
State-of-the-Art: Penwarden et al. (Penwarden et al., 2022) developed multi-fidelity (Penwarden et al., 2022; Lu et al., 2022b; Peherstorfer et al., 2018) neural operators for conjugate heat transfer: Low fidelity: Simplified 1D thermal resistance models (cheap), High fidelity: Coupled CFD-conduction simulations (expensive), and Neural fusion: Multi-fidelity DeepONet learns correction from low to high fidelity
@@ 3.4.1 Parametric Maxwell’s Equations
Time-harmonic Maxwell equations (Chen et al., 2020; Wiecha and Muskens, 2020) in frequency domain:
@@
Application - Metamaterial Design: Chen et al. (Chen et al., 2020) used neural operators for inverse design (Lu et al., 2021b; Meng et al., 2022) of electromagnetic metamaterials.
@@ 3.4.2 Acoustic Wave Propagation
The acoustic wave (Jin et al., 2022a) equation in heterogeneous media:
@@
Seismic Imaging Application: Moseley et al. (Moseley et al., 2020a) addressed full-waveform inversion for subsurface velocity model estimation.
@@ 3.5 Cross-Domain Insights and Application Maturity
Fluid Dynamics Re, Mach, geometry 1-10 FNO, DeepONet, PINO $10^{3}$-$10^{5}\times$ (Li et al., 2020a; Jin et al., 2021)
@@
Industrial Adoption Status: Production deployment: Weather forecasting (FourCastNet (Pathak et al., 2022; Kurth et al., 2023)), some engineering design optimization, Pilot projects: Digital twins, real-time control systems, medical imaging, and Research stage: Nuclear fusion, climate prediction, drug discovery
@@ 4.1.1 The Parametric Solution Manifold
Kolmogorov n-Width: A classical measure of manifold complexity from reduced-order modeling theory (Binev et al., 2011; Berkooz et al., 1993; Hesthaven et al., 2016; Peherstorfer et al., 2014) is the Kolmogorov $n$-width:
@@
For elliptic PDEs with smooth, affine parameter dependence, $d_{n}(\mathcal{M})\sim e^{-\alpha n}$ (exponential decay) justifies reduced basis methods (Binev et al., 2011). However, transport-dominated or hyperbolic problems may have $d_{n}\sim n^{-\beta}$ (polynomial decay), requiring larger $n$ for adequate approximation.
@@ 4.1.2 Universal Approximation Theorems for Operators
Theorem 2 (Neural Operator Universal Approximation (Kovachki et al., 2023b)).
@@
DeepONet: Lu et al. (Lu et al., 2021a) proved that the branch-trunk decomposition equation 28 achieves universal approximation if: Branch network width $p\to\infty$, and Both branch and trunk networks are universal approximators
@@
FNO: Li et al. (Li et al., 2020a) showed Fourier neural operators approximate operators by leveraging spectral properties. For periodic problems, the Fourier series representation provides natural expressivity. The critical assumption: solutions have sufficient spectral decay (smoothness in Fourier space).
@@
where $W$ is network width, $d$ is input dimension, and $\alpha$ is smoothness. The curse of dimensionality ($1/W^{\alpha/d}$) persists theoretically (Han et al., 2018; Beck et al., 2020; Khoo et al., 2021), though empirically neural operators perform better than this bound suggests.
@@ 4.1.3 Parametric PDE-Specific Analysis
Low-Dimensional Structure: Bhattacharya et al. (Bhattacharya et al., 2021) proved that if the solution manifold $\mathcal{M}$ has intrinsic dimension $d_{\text{eff}}\ll d$ (as measured by Kolmogorov width decay), then neural operators can achieve dimension-independent error bounds:
@@ 4.2.2 Sample Complexity Bounds
De Ryck and Mishra (De Ryck et al., 2024b) derived sample complexity for PINNs approximating parametric elliptic PDEs:
@@
For neural operators, Lanthaler et al. (Lanthaler et al., 2022) proved:
@@
Empirical Findings: FNO: Interpolation error $\sim$1-2%, extrapolation error $\sim$10-20% (Li et al. (Li et al., 2020a)), DeepONet: Similar interpolation/extrapolation gap, slightly better extrapolation when physics-informed, and PINN: Poor parametric generalization—essentially no extrapolation capability without retraining
@@ 4.2.4 Out-of-Distribution Detection
Approaches: Uncertainty quantification employs multiple complementary approaches. Bayesian neural operators compute predictive variance as an out-of-distribution indicator. Ensemble methods identify uncertain regions through high variance across ensemble members. Conformal prediction constructs prediction sets with guaranteed coverage (Chalot et al., 2023), providing distribution-free uncertainty bounds.
Breakthrough - Conformal Prediction for PDEs: Staber et al. (Chalot et al., 2023) developed distribution-free uncertainty quantification for neural operators:
@@ 4.4.1 Training Convergence for PINNs
Neural Tangent Kernel (NTK) Analysis: Wang et al. (Wang et al., 2022b), building on failure mode analyses (Krishnapriyan et al., 2021; 2022; Markidis, 2021; Fuks and Tchelepi, 2020), used NTK theory to analyze PINN training dynamics. For infinitely-wide networks in the NTK regime:
@@
This explains empirically observed training failures (Krishnapriyan et al., 2021).
@@ 4.4.2 Operator Learning Convergence
Lanthaler et al. (Lanthaler et al., 2022) proved that for appropriate architecture choices and sufficient training data:
@@ 4.4.3 Long-Time Stability
Recent Success: Brandstetter et al. (Brandstetter et al., 2022) demonstrated stable 1000-step rollouts for turbulent flows by enforcing energy and enstrophy conservation in FNO training.
@@ 5.1.2 Active Subspaces and Dimension Reduction
Many high-dimensional parameter spaces have low effective dimensionality. Active subspace methods (Constantine, 2015) identify important parameter directions.
@@
Case Study: Hu et al. (Hu et al., 2024) addressed a subsurface flow (Tang et al., 2020; Fuks and Tchelepi, 2020) problem with $d=100,000$ uncertain permeability parameters (each grid cell). Active subspace analysis revealed $d_{\text{eff}}=5$ active directions explaining 95% of pressure variance. After reduction: Trained DeepONet on 5D active subspace, Achieved $3\%$ error in pressure predictions, and Total computation: 12 hours on single GPU (vs. years for full-space sampling)
@@ 5.1.4 Latent Variable Models
L-DeepONet: Kontolati et al. (Kontolati et al., 2023) developed latent-space DeepONet for high-dimensional parametric PDEs: Application: Structural dynamics with $d=10,000$ (discretized forcing field) (Rao et al., 2021; Rezaei et al., 2022; Samaniego et al., 2020), Latent dimension: $d_{z}=50$ (learned via variational autoencoder), and Results: 1-2 orders of magnitude speedup vs. standard DeepONet, $5\%$ accuracy loss
@@ 5.2.1 Bayesian Physics-Informed Neural Networks
B-PINNs (Yang et al., 2021) place priors over network weights:
@@
Application - Cardiovascular Flows: Arzani et al. (Arzani et al., 2021), along with related cardiovascular modeling studies (Kissas et al., 2020; Sahli Costabal et al., 2020; Yazdani et al., 2020), used B-PINNs to infer patient-specific blood viscosity and vessel compliance from sparse Doppler ultrasound data. Posterior distributions quantified parameter uncertainty, enabling risk assessment (e.g., probability of flow reversal).
@@ 5.2.2 Bayesian Neural Operators
Efficient Approximations: Laplace approximation: Gaussian approximation around MAP estimate, Concrete dropout: Approximate variational inference via dropout (Gal and Ghahramani, 2016), and Ensemble methods: Train multiple operators with different initializations